OpenStudy (anonymous):

Write using Exponents: (Fourthroot)5t^3 please help!

OpenStudy (lalaly):

\[\sqrt[4]{5t^3}= (5t^3)^{\frac{1}{4}}\]

OpenStudy (anonymous):

\[\sqrt[4]{5t}^{3} \]

OpenStudy (chaise):

(5t^3)^1/4 Multiply the two powers together you get: 5t^(3/4)

OpenStudy (anonymous):

THe answer is apparently supposed to be 5^(1/4)t^(3/4) I just don't know how to get there

OpenStudy (anonymous):

oh then i read it wrong sorry

OpenStudy (lalaly):

\[\sqrt[4]{x}= x^{\frac{1}{4}}\]

OpenStudy (anonymous):

it's just t^3 not the whole radical...

OpenStudy (lalaly):

\[\sqrt[4]{5t^3}= (5t^3)^{\frac{1}{4}} = 5^{\frac{1}{4}} \times t^{\frac{3}{4}}\]

OpenStudy (anonymous):

oh. Thanks :)